### v2

# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xx

Abstract
This paper presents an infrastructure-free approach for obstacle detection and environmental mapping using ultra-wideband (UWB) radar mounted on a mobile robotic platform. Traditional sensing modalities such as visual cameras and Light Detection and Ranging (LiDAR) fail in environments with poor visibility due to darkness, smoke, or reflective surfaces. In these visioned-impaired conditions, UWB radar offers a promising alternative. To this end, this work explores the suitability of robot-mounted UWB radar for environmental mapping in dynamic, anchor-free scenarios. The study investigates how different materials (metal, concrete and plywood) and UWB radio channels (5 and 9) influence the Channel Impulse Response (CIR). Furthermore, a processing pipeline is proposed to achieve reliable mapping of detected obstacles, consisting of 3 steps: (i) target identification (based on CIR peak detection), (ii) filtering (based on peak properties, signal-to-noise score, and phase-difference of arrival), and (iii) clustering (based on distance estimation and angle-of-arrival estimation). The proposed approach successfully reduces noise and multipath effects, resulting in an obstacle detection precision of at least $90.71\%$ and a recall of $88.40\%$ on channel 9 even when detecting low-reflective materials such as concrete. This work offers a foundation for further development of UWB-based localisation and mapping (SLAM) systems that do not rely on visual features and, unlike conventional UWB localisation systems, do not require on fixed anchor nodes for triangulation.
Index Terms:
Autonomous Navigation, Environment Mapping, Obstacle Detection, Radar, SLAM, Ultra-wideband (UWB)
[…]
=-21pt
I Introduction
The ability to detect objects and build a representation of the surrounding environment is an essential requirement to allow autonomous operations of mobile robots. Existing systems rely on cameras to capture images and track movement by extracting visual features. However, this method depends on clear lighting conditions and struggles in environments with low visibility, shadows, or reflective surfaces [review-LiDAR]. Light Detection and Ranging (LiDAR) systems, on the other hand, use laser pulses to measure distances and construct maps. While highly accurate, LiDAR sensors face challenges in environments where laser beams are scattered or absorbed, such as in heavy rain, fog, or areas with multiple reflective surfaces. Given these challenges, there is a growing need for alternative localisation and mapping technologies. Ultra-wideband (UWB) is a promising alternative due to its ability to operate in low-visibility environments while maintaining high accuracy (typically under 20 cm). However, UWB-based localisation systems usually rely on anchor-tag setups, where fixed UWB anchors are installed in predefined locations, and mobile robots are equipped with UWB-tags [ensemble-learning] [nlos-neural-network]. Although this setup works well for environments with stable infrastructure, it is impractical in dynamic or unknown places, such as search and rescue activities or industrial facilities where the surroundings frequently change. Furthermore, using a UWB radio instead of LiDAR can reduce the system’s total weight, a significant advantage for drones that need to be as lightweight as possible to maximize autonomy [survey-robot-swarms].
To realize this vision, this paper proposes and evaluates signal processing methods to detect obstacles and map the environment. For evaluating the proposed algorithm, data was collected using the Qorvo QM33120WDK1 development kit [datasheet-sensors]. The setup consists of separate omnidirectional transmitter antenna and directional receiver antennas mounted on a TurtleBot 4. The main contributions of this paper are as follows:
  • We analyse how three common obstacle materials (metal, concrete, and plywood) affect the Channel Impulse Response (CIR), and identify optimal signal parameters (peak width and prominence) for obstacle detection of these materials across IEEE 802.15.4 UWB channels 5 and 9.
  • We propose a novel noise filtering method that eliminates multipath-induced phantom reflections by computing a reliability score based on peak characteristics, signal-to-noise ratio (SNR) and phase-difference of arrival (PDoA).
  • We broaden the radar’s field of view by clustering detected points using combined distance and angle-of-arrival (AoA) estimations, enabling spatially coherent obstacle mapping.
  • We evaluate the complete processing pipeline in realistic conditions within a large industrial testbed environment.
  • We publicly release the captured dataset, processing pipeline, and an illustrative video to support reproducibility and further research [^1] https://gitlab.ilabt.imec.be/datasets/uwb-radar-obstacle-mapping-dataset.
The remainder of this paper is organised as follows: Section II provides background information on UWB. Section III discusses related work, whereas Section IV presents the proposed processing pipeline. In Section V, the data collection is explained, and the results for the proposed method are presented in Section VI together with a visual demonstration. Finally, Sections VII and VIII cover suggestions for future work and the conclusion, respectively.
II Background of Ultra-Wideband
UWB is a radio technology that transmits data over a wide frequency range, either with an absolute bandwidth greater than 500 MHz or a fractional bandwidth ($B_{F}$) greater than 20%. The fractional bandwidth ($B_{F}$) is defined as the ratio of the signal’s bandwidth ($B_{W}$) to its centre frequency ($f_{C}$). The fractional bandwidth ($B_{F}$) is defined as:
$B_{F}=\frac{BW}{f_{C}}=\frac{(f_{H}-f_{L})}{(f_{H}+f_{L})/2}$
Where $f_{H}$ and $f_{L}$ are the upper and lower frequencies of the -10 dB bandwidth [uwb-overview].
Because UWB technology uses a large bandwidth, it enables the transmission of extremely short, narrow pulses in the time domain as described by the time-bandwidth relation $\text{BW}\times T\geq\frac{1}{\pi}$, where ($B_{W}$) is the signal’s bandwidth and ($T$) the pulse duration. For example, UWB systems using a 500 MHz bandwidth can generate pulses of only 0.16 ns wide, whereas traditional technologies such as Wi-Fi, which are limited to bandwidths around 20 MHz, produce pulses longer than 4 ns. Due to its high bandwidth and ultra-narrow pulses, UWB is considered to have a high range resolution and ability to distinguish multiple targets.
This technology operates within the frequency spectrum of 3.1 GHz to 10.6 GHz, enabling data transmission speeds of up to 110 Mbps, typically with a range of up to 10 meters. A higher range of up to 100 meters is also possible, but at lower data rates, depending on several factors, including channel frequency, antenna design, power levels, and complexity of the propagation environment [uwb-standards].
UWB radar can be implemented either as Direct Sequence UWB (DS-UWB), which spreads a carrier signal using pseudo-random noise code, or as Impulse Radio UWB (IR-UWB), which transmits short pulses and estimates distance using the Time-of-Flight (ToF). Compared to DS-UWB, IR-UWB is simpler to implement, more power-efficient, and more robust against multipath interference, which is why it is most used in radar applications [uwb-overview]. For this reason, this proposed approach focuses on pulse-based UWB.
To better understand how IR-UWB radar performs in realistic environments, it is essential to study how the transmitted signal interacts with obstacles and reflections in the environment. This is described by the Channel Impulse Response (CIR), illustrated in Fig. 1. The first peak in the figure corresponds to the Line-of-Sight (LoS) component, which is the earliest signal arriving directly from the transmitter to the receiver assuming no obstacles in between. In this case, the LoS signal indicates the signal transmitted from the transmitter antenna to the receiver antenna co-located on the same robot. The antennas are separated by a distance of 21 cm. Subsequent peaks represent reflections from other objects, walls, or multipath components.
Figure 1: Example of a Channel Impulse Response where the object is placed at a distance of 3.5 m.
The CIR is typically modelled as a sum of deterministic and diffuse multipath components [cir-formula] as shown in (1):
$CIR(t)=\sum_{i}A_{i}\delta(t-\tau_{i})+\nu(t)$ (1)
The first term represents the deterministic multipath components, which describe the propagation of the UWB signal through multipaths with different delays. Each component is characterized by an amplitude $A_{i}$ and a corresponding delay $\tau_{i}$. The second term accounts for the diffuse multipath components, capturing random variations and scattering effects within the channel. These are modelled as additive white Gaussian noise (AWGN), denoted by $\nu(t)$, to provide a realistic representation of the channel behaviour.
In UWB radar systems, the CIR data is typically obtained as I//Q samples (In-Phase and Quadrature components) of the received signal, which provides important information about the signal’s strength and phase.
UWB radar is widely used in applications that require precise distance measurements, even in cluttered or dynamic environments. Some examples of these applications are presence detection, device-free localisation, monitoring vital signs without physical contact [lambrecht2025lowcostembeddedbreathingrate] and, activity recognition. UWB is a strong candidate for fast-moving robots and dynamic environments due to its low weight, low power consumption, and cost-effectiveness [survey-robot-swarms]. While RGB cameras are also lightweight and energy-efficient, they suffer in complex or low-light environments. LiDAR offers high-accuracy long-range sensing but is heavy and power-intensive. LiDAR performance degrades in environments with smoke, dust, or fire, as laser beams cannot penetrate these particles. Additionally, both LiDAR and depth cameras have difficulties detecting transparent objects. Vanhie-Van Gerwen et al. [indoor-drone] show that across various drone sizes, UWB stands out for its favourable accuracy-to-cost ratio. It also provides precise positioning thanks to its wide bandwidth and short pulses, is resistant to narrowband interference, and respects privacy by not capturing identity-related data. Both mmWave and UWB can distinguish small objects thanks to their high bandwidth. However, signals with lower centre frequencies generally propagate farther. mmWave operates between 30 GHz and 300 GHz, resulting in high-speed data transmission, but the technology suffers from high path loss, limited penetration, and short range. In contrast, UWB experiences less attenuation over distance and better penetration of walls and obstacles.
III Related Work
TABLE I: Overview of Related Work in UWB-Based Object Detection and Distance Estimation
Paper Goal Approach Chip UWB Mobile robot No Anchor Nodes Distance Estimation Angle Estimation of Objects IEEE 802.15.4 Compliant Target Object Detection Accuracy
[cir-processing] Object/person detection using CIR analysis Three CIR processing techniques (ICIR, UCIR, ACIR) with BMA and BSDA algorithms DW1000 - - whiteboard, large tv, metal box, person Avg. error $\leq$ 9 cm using ACIR (larger for persons or smaller objects)
[machine-learning-train] Detect humans/obstacles in automatic train pairing. Supervised Machine Learning P440 from TimeDomain - - - - humans/obstacles Best detection accuracy: 55%-70% indoors,>> 95% outdoors
[distance-estimation-ir-uwb] Remove clutter to improve detection accuracy Low-rank approximation and singular value decomposition methods not specified - - - walking person Avg. error <20<20 cm, RMSE <50<50 cm
[LSTM-based-system-multiple-obstacle-detection] Distinguish objects from noise in transportation use cases. Uses Discrete Wavelet Transform to extract features and forward these to the LSTM network. UMAIN HST-D3 directional antenna - - - - pedestrian, cyclist, vehicle, tram 72.78% precision, 71.34% recall, 72.06% F1-score
[entropy-based] Detect the target object Utilises signal’s entropy UMAIN HST-D3 directional antenna - - - - pedestrian, cyclist, vehicle, tram 75.34% precision, 63.06% recall, 68.65% F1-score
[localisation-navigation-uwb] Signal processing chain for obstacle detection Background subtraction, CFAR and Kalman filter M-sequence UWB - - wall and glass pane not specified
[cooperative-uwb-slam] UWB-SLAM Kalman Filter to reduce followed by particle filter M-Sequence UWB - - walls and corners $\leq$ 20 cm
[anchorless-uwb-1] UWB-SLAM wheel odometry and Extended Kalman Filter X4M300 - - vertical metal rods 6.2 cm localisation error
[anchorless-uwb-2] UWB-SLAM in feature-deficient areas Extends previous approach with UWB AoA anchor-tag for better loop closure X4M300 - - - vertical metal rods 10.3 cm localisation error
This paper Reduce noise and unwanted multipath for reliable obstacle mapping Filtering on peak properties, SNR-score, PDoA and clustering QM33120WDK1 metal plate, concrete and plywood box MAE $\leq$ 11.50 cm, $\geq 90.71\%$ precision, $\geq 88.40\%$ recall on channel 9
UWB is widely known for its use in localisation and anchor-tag positioning, where fixed UWB infrastructure (‘UWB anchor nodes’) is used to localise mobile robots. Although highly accurate, these localisation approaches require significant infrastructure investments and are hence outside the scope of this paper. Instead, in this section, we discuss prior works that focused on obstacle detection using fixed or mobile UWB radar systems and compare how they differ from our work.
Table I shows a comparison of the existing work with our approach. Most state-of-the-art solutions rely on the installation of fixed UWB radar systems. For example, Van Herbruggen et al. [cir-processing] introduce Accumulation Channel Impulse Response (ACIR) technique, which improves signal resolution for obstacle detection. Although this method increases the detection accuracy, it requires the collection of sufficient samples before it can be applied, limiting the maximum speed of robotic platforms. From the research of Sattiraju et al. [machine-learning-train] it is concluded that Random Forest and Extra Trees have the highest accuracy in obstacle detection. However, these methods are prone to overfitting the training data, which makes the model generalise less in new environments. Yun et al. [distance-estimation-ir-uwb] suppress noise from the received signal using singular value decomposition (SVD). Although their method achieves high accuracy, the required matrix computations are computationally intensive and may increase the latency and memory usage for real-time mobile robots. Mimouna et al. [LSTM-based-system-multiple-obstacle-detection], [entropy-based] focus on obstacle detection, but lack characterisation of the obstacle (e.g. estimation of distance and width). Moreover, the use of complex deep learning models in [LSTM-based-system-multiple-obstacle-detection] would lead to a higher execution time. C. Smeenk et al. [localisation-navigation-uwb] present a signal processing chain for M-sequence UWB radars mounted on a mobile robot. Their approach combines background subtraction to remove the antenna crosstalk, with a constant false alarm rate (CFAR) method for object detection and distance estimation. However, the localisation is limited to reporting that an object is in front of the antenna and its distance, without providing its relative position in the environment. The method presented in this work overcomes all the above limitations by using clustering that is more adaptable to unseen and dynamic environments. Moreover, our work also characterizes the position and width of detected obstacles by combining both distance estimation and AoA information.
To the best of our knowledge, only four prior publications focus on robot-mounted UWB radar systems.
[anchorless-uwb-1] and [anchorless-uwb-2] both focus more on Simultaneous Localisation and Mapping (SLAM) localisation aspects than UWB signal processing optimisations for obstacle characterisation. They are also limited by recognizing only vertical rods rather than reconstructing full obstacle representations from various material types. The proposed approach from our work can serve as a complementary method to [anchorless-uwb-1] and [anchorless-uwb-2], aiming to suppress noise and mitigate ghost detections while increasing the overall performance of their proposed UWB-SLAM algorithm. Unlike [anchorless-uwb-2], which complements UWB-SLAM with localisation by employing AoA nodes for robot localisation and loop closure, our paper focuses on estimating the AoA of detected objects using the PDoA between antennas on the same chip. R. Zetik et al. [cooperative-uwb-slam] also focus on SLAM and demonstrate good localisation accuracy across different room shapes. However, their approach relies on proprietary antenna designs tailored for their experiments, which increases deployment cost.
Finally, almost all prior work relies on custom UWB radar-only hardware devices with larger bandwidths and thus higher range resolution. Instead, our proposed approach is fully compliant with low-cost IEEE 802.15.4 standard off-the-shelf hardware. Using this setup, the same hardware can also localise the robot if fixed infrastructure is available, by reusing the transmitted wireless signal for multiple purposes simultaneously. This results in reduced energy consumption, less need for separate hardware, lower costs and weight, and more efficient use of the sparse spectrum. In this way, low-cost joint sensing, communication, and localisation for robotic platforms can be realised [integrated-sensing-and-communication].
IV Proposed approach
Figure 2: Overview of the proposed approach: Each capture contains the preamble CIR and two additional CIRs from both antennas on the same chip used for calculating the PDoA. A new measurement happens every 10.42 ms and contains the data of the environment at one timestamp. These are processed independently followed by filtering out noise or unwanted multipath components using peak properties, SNR-score, and PDoA. The distance to the object is estimated and the AoA is then calculated. Subsequently, multiple rows are combined to cluster reflections from the same object and map the environment.
This section outlines the complete workflow of the proposed approach for environment mapping. As illustrated in Fig. 2, it consists of three main steps: (i) target identification, (ii) signal filtering, and (iii) the combination of the individual observations for obstacle and environment mapping.
IV-A Step 1: Target Identification
UWB offers high time resolution due to its large bandwidth and narrow pulses. This enables the precise differentiation in time of multiple reflected signals. An illustration of this effect is provided in Fig. 3, which shows the CIRs measured for metal, concrete, and plywood obstacles, overlaid on a background CIR measured without any obstacles. The reflections caused by each material appear as distinct peaks at specific delays relative to the first path.
The UWB receiver generates three types of CIRs:
  • Preamble CIR: captured on the first antenna and used for distance estimation.
  • STS1-CIR: obtained from the Scrambled Timestamp Sequence (STS) on the first antenna.
  • STS2-CIR: obtained from the STS after the chip switches to the second antenna.
The QM331x0W chips, which comply with the IEEE 802.15.4z standard, integrate two antennas on the same chip. This standard introduces the STS field mainly for security purposes, but it also enables the creation of additional CIRs that can be used for angle of arrival (AoA) estimation as the same signal reaches each antenna with slightly different phase. On the first antenna, the receiver captures both the preamble CIR and the STS1-based CIR. Then, halfway the UWB packet, the chip switches to the second antenna to capture the STS2-based CIR [single-anchor-localization].
Figure 3: The CIRs measured for metal, concrete and plywood objects, overlaid on a background CIR measured without any obstacles. The reflections caused by each material appear as distinct peaks at specific delays relative to the first path. The metal object (blue) is clearly visible compared to the background.
TABLE II: Mathematical symbols used throughout this paper
Variable Description Unit
$B_{F}$ Fractional bandwidth -
$B_{W}$ Signal’s bandwidth Hz
$F_{C}$ Centre frequency Hz
$F_{H}$ and $F_{L}$ Upper and lower frequency of the -10 dB bandwidth Hz
$CIR(t)$ Channel Impulse Response -
$CIR_{magnitude}$ Magnitude of a complex I/Q sample -
$\phi$ Phase of an I/Q sample radians
$\mathit{SNR\text{-}score}_{i}$ SNR-based score for peak with index i -
$A_{i}$ Amplitude of peak with index i in the CIR -
$A_{noise}$ The RMS of noise floor in the CIR -
$k$ Linear weighting factor for calculating the SNR-score -
$d_{TX-p}$ Distance between transmitter and object cm
$d_{RX-p}$ Distance between receiver and object cm
ToFToF Time of Flight ns
$c$ Speed of light ($3.0\times 10^{8}$m/s)
$d_{TX\text{-}RX}$ Distance between the transmitter and receiver cm
$p$ Point on ellipse indicating the target -
$\alpha$ Phase difference between the two antennas on the same chip radians
$\theta$ Angle of arrival radians
TPTP True positive, the peak that indicates a real object -
FPFP False positive, the peak that indicates a real object, but is actually noise -
FNFN False negative, the peak that indicates a real object, but is filtered out -
$P$ Precision calculated as TP / (TP + FP) -
$R$ Recall calculated as TP / (TP + FN) -
F1-score Calculated as 2 x (P x R) / (P + R) -
IV-A1 Signal processing
To reliably extract the peak indicating an object, every raw CIR undergoes the following processing phases. Firstly, the CIR samples are initially recorded as I/Q components. To calculate the magnitude of the received signals,  (2) is used on the Preamble CIR and  (3) is used to calculate the phase of the received signals on both STS-based CIRs.
$CIR_{magnitude}=\sqrt{I^{2}+Q^{2}}$ (2)
$\phi=\arctan\left(\frac{Q}{I}\right)$ (3)
The first four samples of the CIRs are excluded from further processing, as these samples precede the arrival of the first signal and represent the noise floor. However, these initial samples are used later for calculating the Signal-to-Noise-Ratio (SNR) for filtering purposes. At last, to ensure a consistent threshold and comparability across different measurements, the preamble CIR is normalized using min-max scaling.
IV-A2 Peak detection
Peak detection is performed on the preamble CIR by identifying local maxima through comparison with neighbouring values. Subsequently, peaks are filtered based on their properties, as illustrated in the following step.
IV-B Step 2: Filtering
Figure 4: Definition of prominence, width, and the SNR-based score. Prominence is the vertical distance between a peak and its lowest contour line. In this example, the prominence of the potential object’s peak is measured relative to contour line CL1 rather than CL2, since CL2 has a lower value. The width is calculate at half the prominence, using interpolation to determine the intersection points. The SNR-score evaluates the detected peak relative to the noise floor.
The filtering phase is applied to remove peaks that are likely to correspond to noise or unwanted multipath effects. This filtering process is based on four criteria:
  • prominence: represents how much a peak stands out due to its amplitude and its location relative to other peaks. It is defined as the vertical distance between the peak and its lowest contour line (CL), as illustrated in Fig.4. The lowest CL is determined by searching to the left or right until a higher peak is found, or until the end of the CIR is reached. On each side, the minimum amplitude is taken, and the higher of the two values defines the peak’s lowest CL. For the peak generated by a potential object, the prominence is measured between the peak and CL1. Since CL2 has a lower value, the contour line is CL1.
  • width: the minimum required width of a peak to be considered significant, calculated at half the prominence, with linear interpolation used to locate the intersection points (Fig. 4). Narrow peaks are more likely to represent noise, whereas reflections from physical objects typically produce broader peaks.
  • SNR-scorei: the minimum required SNR-score for a detected peak to be considered a real object. The score is calculated as shown in (4).
    $\text{{SNR-score}}_{i}=20\log_{10}\left(\frac{A_{i}}{A_{\text{noise}}}\right)+k\times delay_{i}$ (4)
    Where $A_{i}$ is the amplitude of the detected peak (at index i) and $A_{noise}$ is calculated as the root mean square (RMS) of the noise floor (Fig. 4). Additionally, to account for signal loss caused by the distance, a linear weighting is applied to the score based on a weighting factor $k$ (0.20 in this case) and time delay $delay_{i}$ between the peak index and the first path index. This adjustment helps detect objects farther away as they naturally have a lower signal amplitude.
  • PDoA: The antenna array’s sensitivity and resolution degrade at wider angles, leading to increased multipath interference and reduced angular accuracy. The calculation of the PDoA, shown in  (5), and adapted from [pdoa-aoa], no longer requires synchronization since both antennas are on the same chip.
    $\alpha=\left((-\phi_{A}+\phi_{B}+\pi)\bmod 2\pi\right)-\pi$ (5)
    Here $\phi_{A}$ and $\phi_{B}$ are the phases measured at antenna A and B.
    As such, peaks that have a corresponding PDoA value outside the range of -2.1325 to 2.1325 radians are discarded as estimates beyond this interval are considered unreliable. These bounds are derived using the inverse of (7), which maps an AoA of -45 to 45 to the corresponding PDoA values.
The exact values of the filtering parameters and how they impact accuracy will be discussed in Section VI-A.
IV-C Step 3: Environment Mapping
The final stage of the proposed method combines the processed observations into a representation of the environment. This process involves three iterative steps: distance estimation, AoA determination and clustering. The distance estimation provides the range of each peak, whereas the AoA is used to position the detected peaks relative to the robot. Finally, peaks originating from the same location are grouped through clustering.
IV-C1 Distance Estimation
For obstacle detection, the peak with the highest amplitude (excluding the first path) that meets all the conditions is selected, as it is most likely to correspond to a real object. This is illustrated in Fig. 3. Each peak beyond the first path arrives later because the reflected signal travels a longer time than the direct path. Therefore, the ToF of the reflected signal is used to measure the total distance between transmitter (TX), object, and receiver (RX).
Fig. 5 illustrates an ellipse $E$ whose foci are the TX antenna and RX antenna. Each point $p$ on the ellipse represents a possible target and is defined such that the sum of the distances $d_{TX-p}$ and $d_{RX-p}$ equals the total path length travelled by the signal between TX, target, and RX, as in (6). The total path length equals the ToF multiplied by the speed of light $c$. To resolve the exact location and distance relative to the robot, the angle of arrival must additionally be determined, as described in (7).
Figure 5: Bistatic setup: The ellipse $E$ is defined by the total path length ($d_{TX-p}+d_{RX-p}$) and the fixed positions of the TX and RX antennas, used as the foci of the ellipse. $d_{TX\text{-}RX}$ is the distance between the TX antenna and RX antenna, $d_{TX-p}$ is the distance between TX antenna and target $p$, $d_{RX-p}$ is the distance between target and RX antenna, and $\theta$ is the AoA .
$E={p\in\mathbb{R}^{2}:d_{TX-p}+d_{RX-p}=ToF\times c}$ (6)
IV-C2 Angle of Arrival
To estimate the AoA, the PDoA between the two antennas on the same chip is used. This is calculated as shown in  (5), followed by  (7), which shows the calculation of the AoA [pdoa-aoa].
$\theta=\frac{1}{0.95}\arcsin\left(\frac{\alpha}{\pi}\right)$ (7)
Tests show that the estimated AoA generally follows the ground truth when the object is in front of the robot. Some deviations appear, especially when approaching the object from the side, but these estimates remain useful for filtering.
The distance to the RX antenna can be calculated using the cosine law on the formed triangle (Fig. 5) and the measured total path length.
$d_{TX-p}^{2}=d_{RX-p}^{2}+d_{TX\text{-}RX}^{2}-2d_{RX-p}d_{TX\text{-}RX}\cos\theta$ (8)
This equation contains two unknown distances, $d_{TX-p}$ and $d_{RX-p}$, and can be transformed to (9).
$d_{RX-p}=\frac{d^{2}-d_{TX\text{-}RX}^{2}}{2(d+d_{TX\text{-}RX}\sin\theta)}$ (9)
Here, $d_{TX\text{-}RX}$ denotes the distance between the TX antenna and RX antenna, $d$ is the total path length ($d_{TX-p}$ + $d_{RX-p}$), and $\theta$ is the calculated AoA. This equation calculates the distance between the object and the RX antenna.
IV-C3 Clustering
By applying the filtering methods from Section IV-B and localising the resulting points relative to the robot using the distance and orientation calculations, a point cloud of potential obstacles is generated. In this final step, a clustering algorithm then removes outliers and groups these points into likely obstacle regions. To this end, the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) is used. DBSCAN is well suited for this work because detected components from real objects tend to appear repeatedly in the same location, forming dense groups of points. By identifying these dense regions, DBSCAN can cluster detected components that originate from the same physical object while simultaneously discarding isolated points as noise. The most important parameter for DBSCAN in this work is epsilon (eps), which defines the maximum distance between two samples for one to be considered in the neighbourhood of the other. The second key parameter is min_samples which is the number of samples in a neighbourhood for a point to be considered as a core point, including the point itself (see Fig. 6). Clusters smaller than these thresholds are discarded as noise. In this paper, eps was set to 20 cm and min_samples to 20 detected peaks. Choosing a smaller value would reduce the number of points labelled as noise, but at the risk of forming incorrect clusters that do not correspond to real objects, which would in turn degrade mapping accuracy. The clustering is performed based on peaks mapped to the Cartesian coordinate system and depends on the number of detected peaks across multiple measurements (these could correspond to different robot positions, although the robot may also remain in the same position (see Fig. 6)).
Since this work aims to establish a foundation for anchor-free UWB-based obstacle mapping that can be integrated into broader SLAM systems, autonomous navigation or real-time obstacle avoidance capabilities are not yet incorporated.
Figure 6: Illustration of the clustering process: Peaks detected across multiple robot positions that form dense regions are grouped into clusters and noise is removed.
V Data Collection
The following data collection experiments were conducted to evaluate the effectiveness of the proposed approach and assess its performance under realistic conditions. The data collection experiments were conducted in the controlled Industrial Internet of Things (IIoT) lab at Ghent University [lab-ugent] focusing on object detection and signal behaviour using UWB radar. All experiments were performed using a TurtleBot 4 robot, equipped with one omnidirectional UWB transmitter on the top and multiple dual-antenna, directional UWB receivers on every side depending on the experiment (Fig. 7a). To reduce the impact caused by the first path signal as the transmitter is placed close to the receiver, an absorber is used (Fig. 7b). The specifications of the radar system are listed in Table III. During data capture, the system alternated between channel 5 and channel 9 for each measurement to assess signal behaviour across different frequency bands. The IIoT lab is equipped with a Motion Capture (MOCAP) system, providing millimetre-level accuracy ground truth data as the robot moves.
Figure 7: a) Overview of the antenna placement on the TurtleBot 4 including the distance between the transmitter and receivers. b) Front view of the TurtleBot 4, equipped with the antennas and absorber.
TABLE III: UWB Radar Specifications - IEEE 802.15.4 compliant
Parameter Value
UWB sensor QM33120WDK1 [datasheet-sensors]
Preamble length 512
Frequency 6.5 GHz and 8 GHz
Channel 5 and 9
Bandwidth 500 MHz
Transmitter specification JL159 Omnidirectional single antenna
Receiver specification JL359 Directional dual antenna
Update rate 96 Hz
Number of I/Q samples 50 I/Q pairs
Figure 8: a) metal plate measuring 25 cm x 28 cm b) concrete box measuring 28 cm x 24 cm c) plywood box measuring 31.5 cm x 26 cm
Figure 9: a) First data collection: the robot moves straight forward to the object. The same movement is done three times, once for every material (metal, concrete, plywood). b) Second data collection: two objects are placed, in the front and parallel to the robot while data is captured using two receivers.
V-A First data collection: different materials
The goal of this data collection was to analyse how UWB radar signals behave depending on the type of material. Three different objects were used in the experiments: a metal plate measuring 25 cm x 28 cm (Fig. 8a), a concrete box measuring 28 cm x 24 cm (Fig. 8b), and a plywood box measuring 31.5 cm x 26 cm (Fig. 8c). These materials were selected for their varying electromagnetic properties: the metal plate is highly reflective, the concrete box partially reflects and absorbs signals, and the plywood box primarily absorbs with minimal reflection. As shown in Fig. 9a, the robot moved in a straight line toward a single object, starting from a distance of 340 cm. For this experiment, only the top transmitter and front receiver were active. Due to the use of different materials, this setup is a good validation of the accuracy of recognizing different material types.
V-B Second data collection: multiple objects
The second data collection aimed to explore the ability to detect objects not placed directly in front of the receiver. In this setup, two white plates made of highly reflective material with dimensions of 116 cm by 55 cm were placed as target objects, one directly in front of the robot and the other parallel to its path to test if the front antenna could detect signals from the side or vice versa (Fig. 9b). The robot started from a distance of 420 cm from the front plate and 60 cm from the parallel plate. The distance to the parallel object increased in 60 cm steps up to 300 cm, while the robot was repositioned at the same start point. Since both objects are highly reflective, this setup creates many multi-path components and phantom peaks and is hence a good validation of the efficiency of the filtering algorithm.
VI Results
This section presents and analyses the results from the proposed processing pipeline. It begins with an overview of the filtering parameters, followed by an analysis of target identification, filtering performance and clustering behaviour. This section concludes with a visual demonstration to validate all subcomponents.
The performance of the proposed approach is evaluated using two types of metrics: (i) detection probability and (ii) distance accuracy (expressed as distance error to indicate how well obstacles are positioned relative to the robot).
Detection probability is expressed as Precision P, Recall R, and F1-score, as defined in  (10)-(12):
$P=\frac{TP}{TP+FP}$ (10)
$R=\frac{TP}{TP+FN}$ (11)
$\text{{F1-score}}=2\times\frac{\text{P}\times\text{R}}{\text{P}+\text{R}}$ (12)
Where TP (true positives) are peaks correctly retained that correspond to actual objects, FP (false positives) wrongly kept noise peaks that do not correspond to real objects (e.g. walls or noise), and FN (false negatives) are peaks incorrectly removed that do correspond to actual objects. For detection probability, a margin of 20 cm is used when comparing estimated distances to ground truth: an estimated peak is considered a true positive if it lies within 20 cm of the ground truth object.
Distance accuracy was evaluated in two different ways. First, after target identification and filtering, the detected peaks were assumed to lie directly in front of the antenna, since no AoA information was available. In this case, distance error was computed as the mean absolute error (MAE) between the ground truth and the estimated distance. The results are discussed in Section VI-B1.
Second, when AoA information was included, the retained peaks could be mapped relative to the robot’s position, after which clustering was applied. The distance error was then calculated as the shortest distance between the ground truth coordinates and the estimated object coordinates in the x/y plane. Finally, the mean of these values was taken. The results are discussed in Section VI-B2.
VI-A Filtering parameters
The three key parameters (peak width, prominence, and SNR-score) were optimised using the dataset obtained from the first data collection (Fig. 9a). A grid search over different parameter values was performed, and each configuration was evaluated using the F1-score and the root mean squared error (RMSE) between the estimated and ground-truth distances. The parameters achieving the highest F1-score and lowest RMSE were selected. To further investigate their behaviour, the relationship between distance and both peak prominence and width was analysed for the three different materials (metal, concrete, plywood). Figures 10a and 10b illustrate the correlations in peak prominence and width as a function of target distance.
Figure 10: Correlation between peak properties and distance: a) Prominence decreases with distance for all materials. b) Width increases as the distance increases for all materials.
As seen in Fig. 10a, peak prominence decreases as the distance increases. At short ranges, the reflected signals retain more energy and produce stronger peaks. With increasing distance, signal attenuation and energy loss cause the peaks to stand out less compared to other peaks or the noise floor. Metal consistently produces the highest peak prominence, which is expected given its high reflectivity. Concrete and plywood, by comparison, show lower prominence because they absorb more energy. At distances lower than 100 cm, concrete cannot be detected with the current setup as the peak prominence and width are too low. As distance grows, both concrete and plywood follow a similar downward trend. This limitation is not exclusive to this material. It can affect any material type depending on its reflectivity. Highly reflective surfaces like metal return more energy producing more prominent peaks even when close to the robot. In contrast, materials such as concrete and plywood absorb more energy, resulting in weaker reflections that are more easily overshadowed by the strong first path signal. When the object is within approximately 60 cm, the reflected peak becomes difficult to distinguish due to the strong first path signal and oversaturation of the receiver [comparative-lidar-radar]. Although an absorber was used between the antennas to mitigate this issue, the results show that the problem persists indicating that antenna design also plays a crucial role. This could be resolved by the design of a dedicated UWB antenna array with better isolation between the individual antenna components [uwb-antenna].
Fig. 10b shows that the peak width increases with distance for all three materials. Close to the robot, the peaks are relatively narrow (around 1-2 ns), but at greater distances they broaden to approximately 3-4 ns. The time delay makes the signal arrive over a longer period, which appears as a broader peak in the CIR.
Finally, width, prominence, and SNR-score were tuned using the F1-score, which balances precision and recall. The optimised values are summarised in Table IV, both for specific materials and overall optimal values across different materials. The remainder of this section uses the overall filtering parameters listed in Table IV. These parameters represent the best configuration for mixed-material environments containing metal, concrete, and plywood. If the material in the environment is known beforehand (i.e., only one material type is present), parameters specific to that material type can be used to further improve detection performance.
TABLE IV: Optimal peak width, peak prominence and SNR-score values for different materials and multiple radio channels.
Channel Material Width Prominence SNR-score
5
Metal 1 0.04 25
Concrete 1 0.05 20
Plywood 1 0.02 15
Overall 1 0.05 20
9
Metal 1 0.05 20
Concrete 2 0.03 10
Plywood 0.10 0.01 10
Overall 0.20 0.03 10
VI-B Target identification, filtering, and mapping accuracy
VI-B1 Impact of different materials after filtering
Figure 11: After step 2: Distance accuracy (errors in cm) for different materials (after filtering on width, prominence, snr-score, and PDoA). Metal is the most reliably detected material in both channel 5 and 9, with an error accuracy lower than 15 cm in 95% of the cases. Concrete and plywood show improved results on channel 9, as the algorithm correctly identifies the correct peak most of the time, with an error accuracy lower than 19 cm in 90% of the cases.
The performance of the algorithm after target identification and filtering (on width, prominence, snr-score, and PDoA) differed across materials and channels, the distance accuracy results are shown in Fig. 11 and the detection probability in Fig. 12. In this step, AoA was not considered, and all detections were assumed in front of the antenna.
Metal: Both channels detected the metallic plate with very high accuracy. On channel 5, the MAE was 6.49 cm with precision and recall above $98\%$. Channel 9 further improved accuracy, reducing the MAE to 5.80 cm and achieving nearly perfect precision and recall.
Concrete: Detection was more challenging for the concrete box. On channel 5, the object was only detected starting at 250 cm, with higher variability (MAE 19.92 cm, SD 36.96 cm) and reduced precision ($81.25\%$). Recall remained high ($93.77\%$), but the large number of outliers increased the error. At close range (¡100 cm), the signal was overshadowed by the first path, and the object was not detected on both channels 5 and 9. In contrast, channel 9 considerably improved performance: detection began at 340 cm, MAE dropped to 10.50 cm, and precision increased to $90.71\%$. Fig. 10a and Fig. 10b illustrate that concrete is overshadowed at close distances because the width and prominence are relatively low.
Plywood: Results on channel 5 were the weakest, with low precision ($79.30\%$) and moderate recall ($68.04\%$). This is primarily due to the low amplitude of the peak, which is often close to the noise floor and difficult to differentiate from background noise. More reliable detection occurred when the robot was within approximately 120 cm. Channel 9 improved performance noticeably: MAE decreased to 7.99 cm (from 30.73 cm), and both precision ($98.14\%$) and recall ($99.23\%$) were near perfect, with consistent detection from 340 cm onward.
Overall, channel 9 consistently outperformed channel 5 for non-metallic materials, resulting in lower MAE, fewer outliers, and higher precision/recall. Fig. 11 shows the higher error of concrete and plywood in channel 5 compared to channel 9. In addition, further accuracy gains can be obtained during step 3 of the processing pipeline (see next sections).
(a) Channel 5
(b) Channel 9
Figure 12: After step 2: Detection performance metrics (precision, recall, F1-score) for each material and channel configuration. Metrics are based on correctly retained object peaks (true positives), wrongly kept noise peaks (false positives), and missed object peaks (false negatives) within the range of 20 cm from the ground truth.
VI-B2 Impact of combined filtering and clustering steps
After target identification and filtering using the different methods, the AoA is calculated. The detected peaks are then mapped relative to the robot, followed by clustering. This step further improves accuracy because it removes additional unwanted peaks that correspond to noise, thereby increasing the detection and distance accuracy.
The evolution of the complete algorithm is illustrated in Fig. 13. For this visualisation, data was collected using the parallel antenna in an environment with two objects (illustrated in Fig. 9b). The mobile robot moves in a straight line from start (indicated with a semi-transparent blue circle) to end (indicated with a fully opaque blue circle). Each sub-figure includes a colour bar on the right-hand side representing the SNR-score where a brighter colour indicates higher SNR-score and thus better signal quality. Table V summarises the key parameters used in this work.
TABLE V: Key parameters used in the proposed approach
Parameter Threshold Comments
width, prominence 1, 0.05 Peak properties: minimum required width and prominence of a peak to be considered significant.
SNR-score 20 Peaks with a SNR-score lower than the threshold are discarded.
AoA_range [45°,45°] Only peaks within this range are kept.
eps 20 cm DBSCAN parameter: max distance between two samples to be part of the same neighbourhood.
min_samples 20 DBSCAN parameter: number of samples in a neighbourhood for a point to be considered as a core point.
min_peaks 50 Minimum number of new peaks before clustering is applied.
range 20 cm Detected peaks within this range are classified as true positives.
bias 15 cm Calibration offset to correct systematic errors in distance measurements.
Figure 13: Overview of the impact of each of the multi-stage processing pipeline steps for object detection. For visual clarity, only the point cloud from the right-side antenna is shown. a) Peak detection: 5601 initial peaks b) Filtering on width, prominence and SNR-score: 444 remaining peaks c) Additional filtering on PDoA: 225 remaining peaks d) Calculating object position relative to the robot using AoA: 225 remaining peaks e) Proposed approach: Peak detection, filtering on peak properties, SNR-score, PDoA, calculating object position and clustering: 89 remaining peaks. The front object is not marked as this is outside the filtered 45° - 45° field of view range of the side antenna.
In the first step (Fig. 13a), a radargram-style visualisation illustrates the robot’s path and the signal quality of the detected peaks, represented by their SNR-score. The data are plotted in an x/y plane to facilitate understanding of subsequent processing steps. In this step no angle of arrival is used, and every detected peak is shown in front of the antenna. At this point, many unwanted peaks caused by noise and unwanted multipath effects are visible, which makes raw CIR unsuitable for direct mapping since it could easily lead to false object detections. To remove the unwanted peaks, the data is filtered in the next stage (Fig. 13b) using width, prominence, and SNR-score defined in Table V. This step significantly reduces unwanted peaks and reveals a clear pattern. It also highlights an important behaviour of the parallel antenna: although it is mounted on the right side of the robot, it still detects peaks created from objects that are not directly in front of it. In this example, peaks from an object located in front of the robot but outside the parallel antenna’s forward-facing range are still recorded, resulting in a line of phantom peaks as shown in the figure diagonally down to the right.
To avoid such misleading detections, the following step (Fig. 13c) introduces filtering based on the PDoA. Only peaks with AoA between 45° - 45° are retained, since research has shown PDoA to be reliable within this range. Beyond these limits, phase wrapping can produce misleading estimates. Once this filtering is applied, the AoA is calculated from the phase difference (Fig. 13d) and the peaks are mapped relative to the robot’s location, which in this case is tracked using the MOCAP system. The final step (Fig. 13e) applies clustering to the remaining peaks and all remaining peaks within a cluster are shown. Detected components that are close to each other are grouped into clusters, which likely represent real objects. The clustering is performed after more than min_peaks peaks have been recorded. Peaks that do not belong to any cluster are considered noise and discarded. As a result, the remaining clusters are assigned consistent SNR-based scores, which strengthens the confidence that they correspond to actual objects in the environment.
The accuracy of the end result of the overall processing pipeline is evaluated in Fig. 14, which presents the cumulative distribution function (CDF) of the errors compared to the ground truth. The results show a median error of 8.48 cm. Moreover, 90% of the errors are below 23.50 cm and 95% below 25.65 cm, demonstrating that the proposed method achieves high accuracy with low error rates.
Figure 14: After step 3: CDF illustrating the errors between the estimated object position and ground truth, 90% of the errors are below 23.50 cm and 95% below 25.65 cm.
VI-B3 Processing latency
Next, this section examines the processing overhead of the full approach. The proposed pipeline achieves an average processing time of 0.48 ms per sample, and 4.73 ms when clustering is applied after 50 samples. The average processing latency is measured on a Windows 11 laptop with an Intel(R) Core(TM) Ultra 7 165U CPU and 16 GB RAM. Both values remain below the 10.42 ms interval between two samples at 96 Hz (Table III).
VI-B4 Visual Demonstration of the Proposed Approach
From the experimental results, it can be concluded that channel 9 achieves the highest performances across all three tested material types. To provide a clearer understanding of this outcome, a visual demonstration is made available online [^2] https://www.youtube.com/watch?v=maDkhShUsjw. The algorithm is not executed in real time in this instance, instead, the recorded data is processed offline and subsequently aligned with the recorded video. Fig. 15 illustrates the experimental setup, which consists of four obstacles (metal, concrete, and plywood), the mobile robot, and its followed path. Fig. 16 presents the resulting environment map generated using our proposed approach.
Figure 15: The lab setup containing four obstacles, the robot and its followed path.
Figure 16: The resulting map showing the detected obstacles and their localisation using our proposed approach.
VII Future Work and Limitations
This research identified two main limitations. First, when the robot approaches the object within approximately 60 cm, the object is no longer detected. This distance threshold was determined experimentally. At distances below 60 cm, the object’s reflection peak becomes too close to the direct-path signal, causing it to be overshadowed. Additionally, multipath components generated by the direct signal may further complicate the separation of the object’s peak. While this effect is primarily due to antenna design and the dominance of the direct path, the object’s material properties also influence the outcome. Highly reflective materials generate stronger peaks, whereas absorptive materials produce weaker responses that are more easily masked. To ensure reliable detection at closer distances, improved antenna isolation between the transmitter and receiver could help reduce the direct-path signal, or higher bandwidth settings could be used to achieve better timing resolution. Increasing the attenuation between the TX antenna and RX antenna can also help reduce direct-path interference, for example by utilizing dedicated UWB antenna arrays. Second, the analysis showed that peak prominence decreases and width increases with distance, however, the filters used in this work employed fixed thresholds. In future work, adaptive thresholding methods could be implemented to account for distance-dependent variations in peak characteristics, thereby improving detection robustness across different ranges and reducing false positives and false negatives.
VIII Conclusion
This paper presented an infrastructure-free UWB radar-based approach for obstacle mapping on mobile robots. The proposed method consists of three steps: (i) target identification (based on CIR peak detection), (ii) filtering (based on peak properties, signal-to-noise score, and phase-difference of arrival), and (iii) clustering (based on distance estimation and angle-of-arrival estimation). The experimental evaluation across channels 5 and 9 and three material types demonstrated that the proposed system achieved high distance-estimation accuracy and detection accuracy, with obstacle localisation errors below 11 cm and precision above $90\%$ on channel 9. The filtering methods and clustering effectively suppressed noise and multipath reflections without requiring any prior background data, making it well suited for dynamic environments. By integrating AoA estimation, the approach enables object localisation and obstacle mapping, providing a strong foundation for use in SLAM and other real-time robotic applications.
[Uncaptioned image] Adelina Giurea received the M.Sc. degree in Information Engineering Technology from Ghent University, Belgium, in September 2025. She is currently pursuing the Ph.D. degree with the IDLab Research Group, focusing on UWB and BLE.
[Uncaptioned image] Ir. Stijn Luchie was born in Veurne in 1999. He received his M.Sc. degree in Computer Science Engineering from Ghent University, Belgium, in September 2022. Shortly thereafter, in October 2022, he started as a researcher with the Department of Information Technology (INTEC) at Ghent University in the IDLab research group. His primary area of scientific interest centres on ultra-wideband technology, with a specific focus on the self-calibration of anchor networks and the applications of UWB radar.
[Uncaptioned image] Dr. ir. Dieter Coppens received the master’s degree in electrical engineering from Ghent University, Belgium, in 2021 and the Ph.D. degree with the IDLab Research Group in 2025. His research interests are wireless networking, indoor localization system based on Ultra-wideband technology and machine learning.
[Uncaptioned image] Prof. dr. ir. Jeroen Hoebeke received the Masters degree in Engineering Computer Science from Ghent University in 2002. In 2007, he obtained a Ph.D. in Engineering Computer Science with his research on adaptive ad hoc routing and Virtual Private Ad Hoc Networks. Current, he is an associate professor in the Internet Technology and Data Science Lab of Ghent University and imec. He is conducting and coordinating research on wireless (IoT) connectivity, embedded communication stacks, deterministic wireless communication and wireless network management. He is author or co-author of more than 200 publications in international journals or conference proceedings.
[Uncaptioned image] Prof. dr. ir. Eli De Poorter is currently a Professor with the IDLab Research Group, Ghent University, and imec. His team performs research on wireless communication technologies such as (indoor) localization solutions, wireless IoT solutions, and machine learning for wireless systems. He performs both fundamental and applied research. For his fundamental research, he is currently the coordinator of several research projects (SBO, FWO, and GOA). He has over 200 publications in international journals or in the proceedings of international conferences. For his applied research, he collaborates with industry partners to transfer research results to industrial applications, and to solve challenging industrial research problems.
[…]